# ITACLIP: Boosting Training-Free Semantic Segmentation with Image, Text, and Architectural Enhancements

### v1

@@ Abstract
Recent advances in foundational Vision Language Models (VLMs) have reshaped the evaluation paradigm in computer vision tasks. These foundational models, especially CLIP, have accelerated research in open-vocabulary computer vision tasks, including Open-Vocabulary Semantic Segmentation (OVSS). Although the initial results are promising, the dense prediction capabilities of VLMs still require further improvement. In this study, we enhance the semantic segmentation performance of CLIP by introducing new modules and modifications: 1) architectural changes in the last layer of ViT and the incorporation of attention maps from the middle layers with the last layer, 2) Image Engineering: applying data augmentations to enrich input image representations, and 3) using Large Language Models (LLMs) to generate definitions and synonyms for each class name to leverage CLIP’s open-vocabulary capabilities. Our training-free method, ITACLIP, outperforms current state-of-the-art approaches on segmentation benchmarks such as COCO-Stuff, COCO-Object, Pascal Context, and Pascal VOC. Our code is available at https://github.com/m-arda-aydn/ITACLIP.
@@ 1 Introduction
In this study, we present ITACLIP (Image-Text-Architectural Enhanced CLIP), our training-free method, which utilizes architectural changes and enriches the input features to improve the segmentation performance. Our approach outperforms the current state-of-the-art methods on COCO-Stuff [8], COCO-Object [39], Pascal Context [46], and Pascal VOC [19]. As illustrated in Fig. 1, ITACLIP produces more accurate segmentation maps than SCLIP [60] and NACLIP [24].
@@ 3.2 Vanilla Approach
Applying a patch-based classification strategy is the most straightforward approach to utilize CLIP for segmentation. We employ CLIP’s image encoder to generate visual features $X^{visual}=[x^{visual}_{\texttt{CLS}},X^{visual}_{patch}]\in\mathbb{R}^{(N+1)\times d}$ of given image $I$, as described in Sec. 3.1. $x^{visual}_{\texttt{CLS}}\in\mathbb{R}^{1\times d}$ and $X^{visual}_{patch}\in\mathbb{R}^{N\times d}$ denote the extracted features from the [CLS] token and image patches, respectively. For patch-based classification, we use only the features obtained from patches, $X^{visual}_{patch}$. CLIP’s text encoder is leveraged to compute the textual embeddings, $X^{test}\in\mathbb{R}^{C\times d}$, for the target $C$ classes. Then, we use cosine similarity to determine similarity scores between the text embeddings and patch features. After the softmax operation, we interpolate the resulting logit to resize it to the original image dimensions. Finally, we perform a class-wise argmax operation to produce the segmentation map $S\in\mathbb{R}^{H\times W}$. Formally,
$S=\operatorname*{argmax}_{c\in C}(upsample(cos(X^{visual}_{patch},X^{test})))$ (5)
@@ 3.4 LLM-based Auxiliary Text Generation
Method Post-process COCO-Stuff COCO-Object VOC Context
Baseline - 7.1 8.6 20.3 9.0
ReCo [55] - 14.8 15.7 25.1 19.9
GroupViT [64] - 15.3 27.5 52.3 18.7
TCL [10] PAMR 19.6 30.4 55.0 30.4
MaskCLIP [69] - 14.6 20.6 38.8 23.2
CLIP-DIY [62] - - 31.0 59.9 -
ClearCLIP [32] - 23.9 33.0 51.8 32.6
SCLIP [60] PAMR 23.9 32.1 61.7 31.5
NACLIP [24] PAMR 25.7 36.2 64.1 35.0
TagCLIP [41] - 18.7 33.5 64.8 -
CaR [57] Dense-CRF - 36.6 67.6 30.5
ITACLIP (Ours) PAMR 27.0 37.7 67.9 37.5
@@ 3.5 Image Engineering
We mainly categorize image augmentations into two categories based on whether they change the spatial structure of the image. Augmentations in the first category maintain the spatial arrangement of the image unchanged, whereas augmentations in the second category alter the spatial order of the image, requiring a reversal of the augmentation to preserve spatial information. We apply 1) Gaussian blur and 2) grayscale augmentations for the first category, while 1) horizontal and 2) vertical flips are utilized for the second category. Thus, four augmented images are obtained for each input image.
$L=\lambda L_{1}+(1-\lambda)L_{2}$ (11)
@@ 4.1 Experimental Setup
Datasets. For a fair comparison with previous studies, we evaluate our method on four common semantic segmentation benchmarks using their official validation sets: COCO-Stuff [8], COCO-Object [39], Pascal Context [46], and Pascal VOC [19]. Specifically, COCO-Stuff comprises 171 classes without an explicit background class. COCO-Object is derived from the COCO-Stuff dataset, combining all stuff classes in COCO-Stuff into a single background class and including 80 thing classes from the original dataset. Pascal Context consists of 59 categories and one background class. Pascal VOC contains 20 object classes in addition to one background class. The validation splits of these datasets include 5000, 5000, 5104, and 1449 images, respectively. We assess the segmentation performance of our method using the mean Intersection over Union (mIoU) metric.
Implementation Details. We use the CLIP-ViT-B/16 [48] model with OpenAI’s pre-trained weights. We define a set of potential background classes for the Pascal VOC and COCO-Object datasets, similar to previous works [60, 57, 41]. This set is fed into CLIP’s text encoder to generate background representations. Additional details about the background set are provided in the Appendix. We utilize ImageNet [12] prompt templates used in CLIP to construct prompt-engineered texts. Following previous studies [60, 57, 24], we apply the text expansion technique that uses additional captions for specific classes (e.g., for “person” class: person, person in shirt, person in dress). Furthermore, we perform slide inference with a 224 × 224 window and a stride of 28, after resizing the short side of images to 336, following the procedure described in [60]. Finally, we employ Pixel-Adaptive Mask Refinement (PAMR) [4] to reduce noise in our predictions.
@@
Method FFN Stuff Object VOC Context
ITACLIP 26.3 36.9 66.3 36.3
ITACLIP 27.0 37.7 67.9 37.5
Table 3: Removing the feed-forward block. “Stuff” and “Object” refer to the COCO-Stuff and COCO-Object, respectively.
@@ 4.2 Main Results
Tab. 1 presents the segmentation performance of ITACLIP compared to other approaches. ITACLIP outperforms current state-of-the-art (SoTA) methods on the COCO-Stuff, COCO-Object, and Pascal Context datasets by a notable margin. In addition, it achieves state-of-the-art performance on Pascal VOC, though with a relatively narrow margin. Furthermore, we emphasize that our method is more robust across all four datasets, whereas most other models experience a performance drop on at least one dataset. For instance, the CLIP-ES-based [40] methods listed in this table—CaR [57] and TagCLIP [41]—fail to maintain their segmentation capability on datasets containing numerous “stuff” classes (e.g., see the results of CaR on Pascal Context and TagCLIP on COCO-Stuff). This performance degradation limits the applicability of these models in real-world scenarios.
@@
Method PAMR Stuff Object VOC Context
ITACLIP 26.3 36.4 65.6 36.0
ITACLIP 27.0 37.7 67.9 37.5
Table 5: Influence of post-processing operation. Comparing the performance of ITACLIP with and without PAMR on all datasets.
@@ 4.3 Ablation Study
Study on selected intermediate layers ($l^{\prime}$). We analyze the impact of various layers to determine which intermediate layers ($l^{\prime}$) yield the best performance; Tab. 4 summarizes our findings. To clearly observe the effect of layers, we evaluate the model without applying PAMR. Based on the visualization in Fig. 3 and a recent study [21] showing that attention maps from the last few layers have a greater effect on the representation, we concentrate more on the middle-deep layers in our analysis. Empirically, employing the 7th, 8th, and 10th layers as intermediate layers results in superior performance. We also report the score using only the final layer’s attention map to demonstrate the effectiveness of our multi-layer approach.
@@
Influence of post-processing operation. Before generating the final predictions, we apply PAMR to mitigate noise in our results. We examine the effect of PAMR on our method across all four datasets, as shown in Tab. 5. Although PAMR improves segmentation performance across all datasets, ITACLIP still achieves competitive results and even surpasses the SoTA on some datasets without PAMR. These results demonstrate that our method is highly robust and performs effectively without relying heavily on post-processing.
Method Stride Stuff Object VOC Context
ITACLIP 224 25.3 36.7 66.1 36.9
ITACLIP 112 26.6 37.4 67.1 37.4
ITACLIP 56 26.9 37.7 67.9 37.5
ITACLIP 28 27.0 37.7 67.9 37.5
Table 7: Role of stride value. We investigate the role of the stride value in our method across all four datasets.
@@
Table 8: Effect of Hyperparameters. $\lambda$ and $\alpha$ denote the Image Engineering and Auxiliary Text Coefficients, respectively. The experiments are conducted without applying PAMR.
Effect of Hyperparameters. Tab. 8 illustrates the effect of Image Engineering ($\lambda$) and Auxiliary Text Coefficients ($\alpha$) on the Pascal Context dataset. We observe minimal performance variance within the range of values listed in the table, highlighting the robustness of our method.
@@ 5 Conclusion
This study introduces ITACLIP, which leverages CLIP’s image-level knowledge for dense prediction tasks without requiring pixel-level annotations or additional training. We propose architectural modifications to CLIP’s visual encoder and introduce the Image Engineering and LLM-based Auxiliary Text Generation modules to expand the input representation. As a result of these modifications and modules, our experiments show that ITACLIP achieves state-of-the-art results across four segmentation datasets. Additionally, the Image Engineering module and LLM-based Text Generation strategy can be seamlessly integrated into a range of computer vision tasks. These modules will assist researchers in further improving prediction quality across various vision tasks.
@@ A.1 Additional Implementation Details
LLM-based Auxiliary Text Generation. As detailed in Sec. 3.4, we leverage LLama 3 [18] to generate auxiliary texts for each class name. We use generated definitions as the auxiliary text type for the COCO-Stuff [8], Pascal VOC [19], and Pascal Context [46] datasets and synonyms for the COCO-Object [39] dataset. Fig. 4(a) and Fig. 4(b) illustrate the procedure for generating definitions and synonyms, respectively. Example definitions and synonyms from the Cambridge Dictionary [9] are utilized to guide LLaMa in producing more precise definitions.
@@
Hyperparameter Stuff Object VOC Context
$\lambda$ 0.75 0.75 0.7 0.75
$\alpha$ 0.2 0.1 0.05 0.15

### v2

@@ Abstract
Recent advances in foundational Vision Language Models (VLMs) have reshaped the evaluation paradigm in computer vision tasks. These foundational models, especially CLIP, have accelerated research in open-vocabulary computer vision tasks, including Open-Vocabulary Semantic Segmentation (OVSS). Although the initial results are promising, the dense prediction capabilities of VLMs still require further improvement. In this study, we enhance the semantic segmentation performance of CLIP by introducing new modules and modifications: 1) architectural changes in the last layer of ViT and the incorporation of attention maps from the middle layers with the last layer, 2) Image Engineering: applying data augmentations to enrich input image representations, and 3) using Large Language Models (LLMs) to generate definitions and synonyms for each class name to leverage CLIP’s open-vocabulary capabilities. Our training-free method, ITACLIP, outperforms current state-of-the-art approaches on five popular segmentation benchmarks. Our code is available at https://github.com/m-arda-aydn/ITACLIP.
@@ 1 Introduction
In this study, we present ITACLIP (Image-Text-Architectural Enhanced CLIP), our training-free method, which utilizes architectural changes and enriches the input features to improve the segmentation performance. Our approach outperforms the current state-of-the-art methods on COCO-Stuff [8], COCO-Object [40], Pascal Context [47], Pascal VOC [20], and Cityscapes [12]. As illustrated in Fig. 1, ITACLIP produces more accurate segmentation maps than SCLIP [61] and NACLIP [25].
@@ 3.2 Vanilla Approach
Applying a patch-based classification strategy is the most straightforward approach to utilize CLIP for segmentation. We employ CLIP’s image encoder to generate visual features $X^{visual}=[x^{visual}_{\texttt{CLS}},X^{visual}_{patch}]\in\mathbb{R}^{(N+1)\times d}$ of given image $I$, as described in Sec. 3.1. $x^{visual}_{\texttt{CLS}}\in\mathbb{R}^{1\times d}$ and $X^{visual}_{patch}\in\mathbb{R}^{N\times d}$ denote the extracted features from the [CLS] token and image patches, respectively. For patch-based classification, we use only the features obtained from patches, $X^{visual}_{patch}$. CLIP’s text encoder is leveraged to compute the textual embeddings, $X^{text}\in\mathbb{R}^{C\times d}$, for the target $C$ classes. Then, we use cosine similarity to determine similarity scores between the text embeddings and patch features. After the softmax operation, we interpolate the resulting logit to resize it to the original image dimensions. Finally, we perform a class-wise argmax operation to produce the segmentation map $S\in\mathbb{R}^{H\times W}$. Formally,
$S=\operatorname*{argmax}_{c\in C}(upsample(cos(X^{visual}_{patch},X^{text})))$ (5)
@@ 3.4 LLM-based Auxiliary Text Generation
Method Post-process COCO-Stuff COCO-Object VOC Context Cityscapes
Baseline - 7.1 8.6 20.3 9.0 6.9
ReCo [56] - 14.8 15.7 25.1 19.9 21.6
GroupViT [65] - 15.3 27.5 52.3 18.7 18.5
TCL [10] PAMR 19.6 30.4 55.0 30.4 23.5
MaskCLIP [70] - 14.6 20.6 38.8 23.2 24.9
CLIP-DIY [63] - - 31.0 59.9 - -
ClearCLIP [33] - 23.9 33.0 51.8 32.6 30.0
SCLIP [61] PAMR 23.9 32.1 61.7 31.5 34.1
NACLIP [25] PAMR 25.7 36.2 64.1 35.0 38.3
TagCLIP [42] - 18.7 33.5 64.8 - -
CaR [58] Dense-CRF - 36.6 67.6 30.5 -
ITACLIP (Ours) PAMR 27.0 37.7 67.9 37.5 40.2
@@ 3.5 Image Engineering
We categorize image augmentations into two groups based on whether they change the spatial structure of the image. Augmentations in the first category maintain the spatial arrangement of the image unchanged, whereas second category augmentations alter the spatial order of the image, requiring a reversal of the augmentation to preserve spatial information. We apply 1) Gaussian blur and 2) grayscale augmentations for the first category, while 1) horizontal and 2) vertical flips are utilized for the second category. Thus, each input image generates four augmentations.
@@ 4.1 Experimental Setup
Datasets. For a fair comparison with previous studies, we evaluate our method on five common semantic segmentation benchmarks using their official validation sets: COCO-Stuff [8], COCO-Object [40], Pascal Context [47], Pascal VOC [20], and Cityscapes [12]. Specifically, COCO-Stuff comprises 171 classes without an explicit background class. COCO-Object is derived from the COCO-Stuff dataset, combining all stuff classes in COCO-Stuff into a single background class and including 80 thing classes from the original dataset. Pascal Context consists of 59 categories and one background class. Pascal VOC contains 20 object classes in addition to one background class. The Cityscapes dataset focuses on urban street scenes, with the validation set containing 19 semantic categories and no separate background class. The validation splits of these datasets include 5000, 5000, 5104, 1449, and 500 images, respectively. We assess the segmentation performance of our method using the mean Intersection over Union (mIoU) metric.
FFN Stuff Object VOC Context City
26.3 36.9 66.3 36.3 39.4
27.0 37.7 67.9 37.5 40.2
Table 3: Removing the feed-forward block. “Stuff”, “Object”, and “City” refer to COCO-Stuff, COCO-Object, and Cityscapes.
Implementation Details. We use the CLIP-ViT-B/16 [49] model with OpenAI’s pre-trained weights. We define a set of potential background classes for the Pascal VOC and COCO-Object datasets, similar to previous works [61, 58, 42]. This set is fed into CLIP’s text encoder to generate background representations. Additional details about the background set are provided in the Appendix. We utilize ImageNet [13] prompt templates used in CLIP to construct prompt-engineered texts. Following previous studies [61, 58, 25], we apply the text expansion technique that uses additional captions for specific classes (e.g., for “person” class: person, person in shirt, person in dress). Furthermore, we perform slide inference with a 224 × 224 window and a stride of 28, after resizing the short side of images to 336 (560 for Cityscapes), following the procedure described in [61]. Finally, we employ Pixel-Adaptive Mask Refinement (PAMR) [4] to reduce noise in our predictions.
@@ 4.2 Main Results
Tab. 1 presents the segmentation performance of ITACLIP compared to other approaches. ITACLIP outperforms current state-of-the-art (SoTA) methods on the COCO-Stuff, COCO-Object, Pascal Context, and Cityscapes datasets by a notable margin. In addition, it achieves state-of-the-art performance on Pascal VOC, though with a relatively narrow margin. Furthermore, we emphasize that our method is more robust across all five datasets, whereas most other models experience a performance drop on at least one dataset. For instance, the CLIP-ES-based [41] methods listed in this table—CaR [58] and TagCLIP [42]—fail to maintain their segmentation capability on datasets containing numerous “stuff” classes (e.g., see the results of CaR on Pascal Context and TagCLIP on COCO-Stuff). This performance degradation limits the applicability of these models in real-world scenarios.
@@
PAMR Stuff Object VOC Context City
26.3 36.4 65.6 36.0 39.2
27.0 37.7 67.9 37.5 40.2
Table 5: Effect of post-processing operation. Comparing the performance of ITACLIP with and without PAMR on all datasets.
@@ 4.3 Ablation Study
Study on selected intermediate layers ($l^{\prime}$). We analyze the impact of various layers to determine which intermediate layers yield the best performance; Tab. 4 summarizes our findings. To clearly observe the effect of layers, we evaluate the model without PAMR. Based on the visualization in Fig. 3 and a recent study [22] showing that attention maps from the last few layers have a greater effect on the representation, we concentrate more on the middle-deep layers in our analysis. Empirically, employing the 7th, 8th, and 10th layers as intermediate layers results in superior performance. We also report the score using only the final layer’s attention map to demonstrate the effectiveness of our multi-layer approach.
@@
Effect of post-processing operation. Before generating the final predictions, we apply PAMR to mitigate noise in our results. We examine the effect of PAMR on our method across all datasets, as shown in Tab. 5. Although PAMR improves segmentation performance across all datasets, ITACLIP still achieves competitive results and even surpasses the SoTA on some datasets without PAMR. These results demonstrate that our method is highly robust and performs well without relying heavily on post-processing.
Stride Stuff Object VOC Context City
224 25.3 36.7 66.1 36.9 37.7
112 26.6 37.4 67.1 37.4 39.7
56 26.9 37.7 67.9 37.5 40.2
28 27.0 37.7 67.9 37.5 40.2
Table 7: Role of stride value. We investigate the role of the stride value in our method across all five datasets.
@@
Table 8: Influence of Hyperparameters. $\lambda$ and $\alpha$ denote the Image Engineering and Auxiliary Text Coefficients, respectively. The experiments are conducted without applying PAMR.
Influence of Hyperparameters. Tab. 8 illustrates the effect of Image Engineering ($\lambda$) and Auxiliary Text Coefficients ($\alpha$) on the Pascal Context dataset. We observe minimal performance variance within the range of values listed in the table, highlighting the robustness of our method.
@@ 5 Conclusion
This study introduces ITACLIP, which leverages CLIP’s image-level knowledge for dense prediction tasks without requiring pixel-level annotations or additional training. We propose architectural modifications to CLIP’s visual encoder and introduce the Image Engineering and Auxiliary Text Generation modules to expand the input representation. As a result of these modifications and modules, our experiments show that ITACLIP achieves SoTA results across five segmentation datasets. Additionally, the Image Engineering module and LLM-based Text Generation strategy can be seamlessly integrated into a range of computer vision tasks. These modules will assist researchers in further improving prediction quality across various vision tasks.
@@ A.1 Additional Implementation Details
LLM-based Auxiliary Text Generation. As detailed in Sec. 3.4, we leverage LLama 3 [19] to generate auxiliary texts for each class name. We use generated definitions as the auxiliary text type for the COCO-Stuff [8], Pascal VOC [20], and Pascal Context [47] datasets and synonyms for the COCO-Object [40] and Cityscapes [12] datasets. Fig. 4(a) and Fig. 4(b) illustrate the procedure for generating definitions and synonyms, respectively. Example definitions and synonyms from the Cambridge Dictionary [9] are utilized to guide LLaMa in producing more precise definitions.
@@
Hyperparameter Stuff Object VOC Context City
$\lambda$ 0.75 0.75 0.7 0.75 0.7
$\alpha$ 0.2 0.1 0.05 0.15 0.05